OpenStudy (anonymous):

giving medal for answer What is the measure of each interior angle of a regular heptagon? Round to the nearest tenth. a. 96.2° c. 112.2° b. 98.6° d. 128.6°

OpenStudy (anonymous):

The equation for the sum of the angles in a polygon of n sides is: \[180(n-2)\]Since it is a 7 sided figure, we plug in n=7: \[180(7-2)=180 \times 5 = 900 degrees\]Since the sum of the seven angles is 900, each angle will be: \[\frac{ 900 }{ 7 } degrees\]

OpenStudy (anonymous):

0.0